# Generalized Out-of-Distribution Detection and Beyond in Vision Language Model Era: A Survey

### v1

@@ 5 CLIP-based AD: Methodology
In this section, we introduce methodologies for CLIP-based anomaly detection (AD) in the hope that the contrast with OOD detection clarifies the similarities and differences between each task and facilitates a deeper understanding of CLIP-based OOD detection.
@@ 5.1 Zero-shot Anomaly Detection
CLIP-based zero-shot AD was proposed in 2023 by Jeong et al. [27]. Although it started about two years later than OOD detection, many methods have been proposed up to the present.
Definition of Zero-shot AD  The meaning of the term “Zero-shot” for zero-shot AD is similar to that for zero-shot OOD detection. In zero-shot AD, the term “Zero-shot” refers to the non-use of the images in the target domain during both training and inference phases. For instance, the method with additional training with auxiliary datasets can be regarded as a zero-shot method [63, 64, 62, 29, 65]. The method with the pre-processing of the target class texts can also be regarded as a zero-shot method [27, 61, 62].
@@ 5.1.1 Training-free Methods
With Anomaly Prompts  In zero-shot AD, a common approach is to utilize anomaly prompts to detect anomalies. This hypothesis is supported by several observations from existing work [27]. Firstly, the concepts of normality and anomalies are context-dependent states [125] of an object, with language playing a crucial role in defining these states. Secondly, language provides additional insights that help differentiate defects from acceptable variations in normality.
The simplest zero-shot AD methods are (i) to perform anomaly classification with CLIP using text prompts for normality and anomalies as classes (i.e., “normal [class]vs. “anomalous [class]”) and (ii) to calculate the similarity to the normal prompt (i.e., “normal [class]”) as the score. These methods are called CLIP-AC [27]. Jeong et al. [27] reported that CLIP-AC with both normal and anomaly prompts outperforms that with only normal text prompts, which indicates the importance of the use of anomaly prompts. However, the performances for this naive method are not yet satisfactory due to the wide range of variations of anomalies. To solve this issue, Jeong et al. [27] proposed WinCLIP. WinCLIP performs a compositional ensemble on a large number of pre-defined normal and anomaly templates and efficient extraction and aggregation of window/patch/image-level features aligned with the text. WinCLIP outperforms CLIP-AC by a large margin. Because of its simplicity and pioneering work, WinCLIP has become an important baseline for CLIP-based AD. AnoCLIP [61] follows WinCLIP’s approach of using a large number of pre-defined normal and anomaly templates but modifies the templates to be domain-aware (e.g., industrial photo) and contrastive state for normal and anomaly (e.g., perfect and imperfect). However, it is noteworthy that the performance of the ensemble strategies of previous methods heavily depends on the text descriptions [27, 61]. Also, it is observed that more descriptions are not always better [62], which makes the previous approaches [27, 61] using a naive ensemble of large templates somewhat uncontrollable and random in their applications. Therefore, SDP [62] proposes RVS, a representative vector selection paradigm, which makes the mechanism of extracting representative vectors from large templates controllable, allowing for a more diverse selection of representative vectors.
@@ 5.1.2 Auxiliary Training-based Methods
Unlike CLIP-based OOD detection, all auxiliary training-based methods in zero-shot AD are open-vocabulary AD methods that are trained with auxiliary AD datasets and tested on unseen target datasets by simply changing the category prompt. Existing methods perform supervised training on the test set of one dataset and perform zero-shot testing on the other dataset [63, 29, 65] (e.g., training with MVTec-AD and evaluation with VisA.)
@@
In recent years, the development of auxiliary training-based methods has received more attention than training-free zero-shot methods. There are two main reasons why training is necessary for AD: (i) The first is the domain gap between semantics and anomalies. CLIP is pre-trained to understand the semantics of images, so when applied in a zero-shot manner, it captures the semantics of the image. However, actual anomalies are not semantics, but rather represent the state of an object and appear only in local areas of the image. Therefore, without training, this domain gap between semantics and anomalies cannot be bridged. (ii) The second reason is that there are limitations to relying on a large set of manually crafted anomaly prompts. This incurs prompt creation costs and also makes it difficult to respond to unknown anomalies. Therefore, by replacing the anomaly prompts with learnable parameters, they aim to solve the high costs and limited adaptability to new anomalies.
@@
To address the above issue (i), APRIL-GAN (also known as VAND) [63] was proposed. APRIL-GAN tackles the domain gap between semantics and anomaly by adding additional linear layers in vision encoders. These linear layers project image features at each scale into the text space, creating and aggregating anomaly maps at each stage. Similarly, SDP+ [99] also incorporates additional linear layers into SDP [99] to effectively project image features into the text feature space, addressing the misalignment between image and text. To solve both the issue (i) and (ii), AnomalyCLIP [29] was proposed. AnomalyCLIP is a prompt learning-based method similar to CoOp. By replacing anomaly prompts with learnable parameters, it eliminates the need to prepare a large number of manually pre-defined prompts such as those in WinCLIP [27]. Furthermore, unlike CoOp which learns object semantics, AnomalyCLIP learns object-agnostic text prompts that capture generic normality and abnormality in an image regardless of its semantics. To achieve this, AnomalyCLIP introduces object-agnostic text prompt templates for both normal and anomaly and performs global and local context optimization. A more recent approach, FiLo [65] leverages Large Language Models (LLMs) to generate fine-grained anomaly descriptions for each object category. This method replaces generic abnormal descriptions with LLM-generated specific anomaly content for each sample. By adding learnable prompts before the generated anomaly prompts, FiLo performs global and local context optimization, enhancing the ability to detect anomalies. As a unique direction from these methods, RWDA [64] proposes a data augmentation approach by utilizing CLIP’s text embeddings as training data. RWDA adds randomly generated words into normal and anomaly prompts to generate a diverse set of normal and anomaly training samples and trains a regular feed-forward neural network with diverse text embeddings.
@@ 5.2 Few-shot Anomaly Detection
CLIP-based few-shot AD was proposed in 2023 by Jeong et al. [27], concurrently with the development of zero-shot AD [27]. Traditional few-shot AD research focuses on modeling the normal distribution from a limited number of normal samples to detect anomalies [126, 127, 128, 129, 130, 131, 132, 133]. However, these methods often struggle to generalize to new domains, as they typically require retraining with the target datasets. With the advent of CLIP, the field of few-shot AD is shifting towards using a few target images only for inference at test time without training.
Definition of Few-shot AD  CLIP-based few-shot AD aims to detect anomaly images using only a few images in the target domain. The meaning of the term “Few-shot” is similar to that of few-shot OOD detection. In few-shot AD, the term “Few-shot” refers to the use of a few normal images in the target domain during training or inference phases. For instance, the method with additional training with a few normal images in the target domain can be regarded as a few-shot method [28, 41]. Even without training, if a method uses a few normal images in the target domain as a reference, we regard it as a few-shot method [27].
@@ 5.2.1 Training-free Methods
The earliest approach in CLIP-based few-shot AD is WinCLIP+ [27], an improved method of WinCLIP. WinCLIP, a base zero-shot AD method, cannot identify certain defects that can only be defined visually rather than textually. For example, the “Metal-nut” category in MVTecAD has an anomaly type labeled “flipped upside down,” which can only be identified relative to a normal image. To address this, WinCLIP+ incorporates a few normal reference images into a memory bank [134] and calculates the anomaly score with the cosine similarity between the query image and its most similar image in the memory bank.
@@ 5.2.2 ID Training-based Methods
There are few training-based methods for CLIP-based few-shot AD research, although CLIP-based few-shot OOD detection actively explores the training-based methods. This is likely because detecting anomalies in unknown classes is highly valuable in practical applications, and training on target data may oversimplify the task of few-shot anomaly detection. This oversimplification concern arises because the anomaly space for known categories is much more limited compared to that for OOD detection. Training with the target data could therefore make the task oversimplified, reducing its difficulty. The only existing work in this area is PromptAD [33], a prompt learning method for one-class AD (where the normal class consists of one class). In one-class AD, traditional prompt learning methods for multi-class classification (e.g., CoOp [24]) do not work well. To address this, PromptAD creates a large number of anomaly prompts by adding a learnable anomaly suffix to the normal prompt. It then learns to bring the visual features closer to the normal prompt and further away from the anomaly prompts, enabling prompt learning for one-class AD.
@@ 5.2.3 Auxiliary Training- and Reference-based Methods
We explore the methods trained on auxiliary datasets and utilize the normal images in the target domain as references during inference. An early work in this category is APRIL-GAN (few-shot) [63], which uses a linear layer trained with auxiliary datasets. Similar to WinCLIP+ [27], APRIL-GAN (few-shot) utilizes a few ID reference images with a memory bank-based approach [134]. More recently, Zhu et al. [35] proposed an in-context-learning-based method called InCTRL. InCTRL trains a model to discriminate anomalies from normal samples by learning to identify residuals or discrepancies between query images and a set of few-shot normal images (in-context sample prompts) from auxiliary data. During inference, InCTRL identifies anomalies by measuring the discrepancy between the features of the query image and a few in-context normal samples from the target dataset.
@@ 5.3.1 Anomaly Detection with Localization Models
Some works [135, 136] tackle AD using SAM [57] or DINO [56], foundation models for localization. The representative work is SAA and SAA+ [135]. SAA is a simple baseline approach and utilizes Grounding-DINO [56] for the anomaly region generator and SAM [57] for the anomaly region refiner. SAA+ is an improved method of SAA and it incorporates domain expert knowledge and target image context into SAA. Given that AD necessitates localization, it is expected that the number of works employing localization foundation models such as SAM will continue to increase.
@@ 5.3.2 Medical Anomaly Detection
While most works on CLIP-based AD focus on industrial AD, recent studies have begun to challenge medical anomaly detection (medical AD) [62, 35, 36, 137, 138]. CLIP-based medical AD is a more challenging area than industrial AD due to the larger gap between different data modalities. A representative work on medical AD is MVFA [36]. MVFA is a method specifically tailored for medical AD. It incorporates multiple residual adapters into the CLIP’s visual encoder to reduce the domain gap, enabling a stepwise enhancement of visual features across different levels. The future progression of medical AD and industrial AD offers an intriguing perspective, exploring whether these fields will develop independently or influence each other. However, when considering practical applications, it should be noted that medical AD faces the challenge that anomalies are not always describable in the language. Therefore, the development of medical AD methods that do not use CLIP is also important.
@@ 5.4 Discussion
We discuss the similarities and differences between CLIP-based OOD detection and CLIP-based AD to deepen our understanding of CLIP-based OOD detection.
@@ 5.4.1 Difference between Each Methodology
Differing Scopes of OOD  OOD detection and AD differ significantly in the scope of OOD (anomaly) they cover, which leads to differences in methodologies, particularly in the use of OOD prompts. As explained in Sec. 3, sensory AD is intended for specific use cases like industrial inspection, where samples deviating from predefined normality (e.g., defective products) are considered anomalies [22, 11]. In other words, in sensory AD, the anomaly space is limited to damaged objects with shared semantics, and anomalies like images of dogs are not expected. This limited anomaly space allows even simple prompts to achieve decent performance. Therefore, as shown in Table II, all AD methods utilize anomaly prompts. Conversely, in OOD detection, as explained in Sec. 3.1, anything semantically different from the ID class is considered OOD. Thus, utilizing naive manual OOD prompts is forbidden (even if it improves benchmark performance). This vastness of the OOD space is the key factor differentiating the methodologies between the two fields.
Essential Features Learned for Detecting OOD  There is a significant difference in the features that need to be captured between CLIP-based OOD detection and CLIP-based AD. In CLIP-based OOD detection, it is ideal to learn a more compact ID decision boundary that produces low uncertainty for the ID data, with high uncertainty for OOD data [28]. On the other hand, CLIP-based AD aims to learn the anomalies, instead of ID decision boundary for detecting anomalies [29]. The learned features are entirely different, which is helpful in developing each training-based method.
Difficulty of Localization Task  CLIP-based OOD detection and CLIP-based AD differ significantly in the difficulty of OOD (anomaly) localization tasks. In CLIP-based AD, anomaly segmentation is a mainstream task, often performed alongside classification in many papers. However, in CLIP-based OOD detection, there has been no research on object-level OOD detection/segmentation. Object-level OOD detection aims to detect OOD objects [139, 140, 141]. The inactivity is related to the size of the OOD space, and the too-vast space of OOD makes it difficult to identify OOD objects with prompts effectively. To pave the way for future development, the foundation models for localization such as SAM [57], which can segment individual objects, have the potential to address object-level OOD detection/segmentation. Object-level OOD detection/segmentation using SAM is a promising future research direction.

### v2

@@ 5 VLM-based AD: Methodology
In this section, we introduce methodologies for VLM-based anomaly detection (AD) in the hope that the contrast with OOD detection clarifies the similarities and differences between each task and facilitates a deeper understanding of VLM-based OOD detection.
Common Findings in All Settings  In VLM-based AD, all representative methods utilize anomaly prompts (e.g., “anomalous [class]”) to detect anomalies. This hypothesis is supported by several observations from existing work (Jeong et al., 2023). Firstly, the concepts of normality and anomalies are context-dependent states (Isola et al., 2015) of an object, with language playing a crucial role in defining these states. Secondly, language provides additional insights that help differentiate defects from acceptable variations in normality.
@@ 5.1 Zero-shot Anomaly Detection
VLM-based zero-shot AD was proposed in 2023 by Jeong et al. (2023). Although it started about two years later than OOD detection, many methods have been proposed up to the present.
Definition of Zero-shot AD  The meaning of the term “Zero-shot” for zero-shot AD is similar to that for zero-shot OOD detection. In zero-shot AD, the term “Zero-shot” refers to the non-use of the images in the target domain during both training and inference phases. For instance, the method with additional training with auxiliary datasets can be regarded as a zero-shot method (Chen et al., 2023b; Tamura, 2023; Chen et al., 2023c; Zhou et al., 2024a; Gu et al., 2024a). The method with the pre-processing of the target class texts can also be regarded as a zero-shot method (Jeong et al., 2023; Deng et al., 2023; Chen et al., 2023c).
@@ 5.1.1 Training-free Methods
The simplest zero-shot AD methods are (i) to perform anomaly classification with CLIP using text prompts for normality and anomalies as classes (i.e., “normal [class]vs. “anomalous [class]”) and (ii) to calculate the similarity to the normal prompt (i.e., “normal [class]”) as the score. These methods are called CLIP-AC (Jeong et al., 2023). Jeong et al. (2023) reported that CLIP-AC with both normal and anomaly prompts outperforms that with only normal text prompts, which indicates the importance of the use of anomaly prompts. However, the performances for this naive method are not yet satisfactory due to the wide range of variations of anomalies. To solve this issue, Jeong et al. (2023) proposed WinCLIP. WinCLIP performs a compositional ensemble on a large number of pre-defined normal and anomaly templates and efficient extraction and aggregation of window/patch/image-level features aligned with the text. WinCLIP outperforms CLIP-AC by a large margin. Because of its simplicity and pioneering work, WinCLIP has become an important baseline for VLM-based AD. AnoCLIP (Deng et al., 2023) follows WinCLIP’s approach of using a large number of pre-defined normal and anomaly templates but modifies the templates to be domain-aware (e.g., industrial photo) and contrastive state for normal and anomaly (e.g., perfect and imperfect). However, it is noteworthy that the performance of the ensemble strategies of previous methods heavily depends on the text descriptions (Jeong et al., 2023; Deng et al., 2023). Also, it is observed that more descriptions are not always better (Chen et al., 2023c), which makes the previous approaches (Jeong et al., 2023; Deng et al., 2023) using a naive ensemble of large templates somewhat uncontrollable and random in their applications. Therefore, SDP (Chen et al., 2023c) proposes RVS, a representative vector selection paradigm, which makes the mechanism of extracting representative vectors from large templates controllable, allowing for a more diverse selection of representative vectors. As a more recent work, ALFA is a zero-shot anomaly detection framework that leverages LLMs to generate adaptive anomaly descriptions at runtime and optimizes the image-text alignment of CLIP at both global and local levels, enhancing detection accuracy and interpretability (Zhu et al., 2024a).
@@ 5.1.2 Auxiliary Training-based Methods
Existing methods with auxiliary training perform training on the test set of one dataset and perform zero-shot testing on the other dataset (Chen et al., 2023b; Zhou et al., 2024a; Gu et al., 2024a) (e.g., training with MVTec-AD and evaluation with VisA.) In recent years, the development of auxiliary training-based methods has received more attention than training-free zero-shot methods. There are two main reasons why training is necessary for AD: (i) The first is the domain gap between semantics and anomalies. CLIP is pre-trained to understand the semantics of images, so when applied in a zero-shot manner, it captures the semantics of the image. However, actual anomalies are not semantics, but rather represent the state of an object and appear only in local areas of the image. Therefore, without training, this domain gap between semantics and anomalies cannot be bridged. (ii) The second reason is that there are limitations to relying on a large set of manually crafted anomaly prompts. This incurs prompt creation costs and also makes it difficult to respond to unknown anomalies.
@@
To address the above issue (i), APRIL-GAN (also known as VAND) (Chen et al., 2023b) was proposed. APRIL-GAN tackles the domain gap between semantics and anomaly by adding additional linear layers in vision encoders. These linear layers project image features at each scale into the text space, creating and aggregating anomaly maps at each stage. Similarly, SDP+ (Chen et al., 2024b) also incorporates additional linear layers into SDP (Chen et al., 2024b) to effectively project image features into the text feature space, addressing the misalignment between image and text. To solve both the issue (i) and (ii), AnomalyCLIP (Zhou et al., 2024a) was proposed. AnomalyCLIP is a textual prompt learning-based method similar to CoOp. By replacing anomaly prompts with learnable parameters, it eliminates the need to prepare a large number of manually pre-defined prompts such as those in WinCLIP (Jeong et al., 2023). Furthermore, unlike CoOp (Zhou et al., 2022c) which learns object semantics, AnomalyCLIP learns object-agnostic text prompts that capture generic normality and abnormality in an image regardless of its semantics. To achieve this, AnomalyCLIP introduces object-agnostic text prompt templates for both normal and anomaly and performs global and local context optimization. FiLo (Gu et al., 2024a) leverages Large Language Models (LLMs) to generate fine-grained anomaly descriptions for each object category. This method replaces generic abnormal descriptions with LLM-generated specific anomaly content for each sample. By adding learnable prompts before the generated anomaly prompts, FiLo performs global and local context optimization, enhancing the ability to detect anomalies. As a more recent method, AdaCLIP (Cao et al., 2024c) employs static and dynamic learnable prompts in both the image and text encoders. Static prompts provide a general adaptation for zero-shot detection, while dynamic prompts are generated per test image for adaptive refinement. Their combination, termed hybrid prompts, enhances zero-shot performance.
As a unique direction from these methods, RWDA (Tamura, 2023) proposes a data augmentation approach by utilizing CLIP’s text embeddings as training data. RWDA adds randomly generated words into normal and anomaly prompts to generate a diverse set of normal and anomaly training samples and trains a regular feed-forward neural network with diverse text embeddings. VCP-CLIP (Qu et al., 2024) proposes to leverage visual context prompting (VCP) to enhance CLIP’s ability to detect and segment anomalies without requiring product-specific text prompts. VCP eliminates reliance on predefined textual descriptions and improves generalization across unseen product categories.
@@ 5.2 Few-shot Anomaly Detection
VLM-based few-shot AD was proposed in 2023 by Jeong et al. (2023), concurrently with the development of zero-shot AD (Jeong et al., 2023).
Definition of Few-shot AD  VLM-based few-shot AD aims to detect anomaly images using only a few images in the target domain. The meaning of the term “Few-shot” is similar to that of few-shot OOD detection. In few-shot AD, the term “Few-shot” refers to the use of a few images in the target domain during training or inference phases. For instance, the method with additional training with a few images in the target domain can be regarded as a few-shot method (Li et al., 2024c). Even without training, if a method uses a few images in the target domain as a reference, it is regarded as a few-shot method (Jeong et al., 2023).
@@ 5.2.1 Training-free Methods
The earliest approach in VLM-based few-shot AD is WinCLIP+ (Jeong et al., 2023), an improved method of WinCLIP. WinCLIP, a base zero-shot AD method, cannot identify certain defects that can only be defined visually rather than textually. For example, the “Metal-nut” category in MVTecAD has an anomaly type labeled “flipped upside down,” which can only be identified relative to a normal image. To address this, WinCLIP+ incorporates a few normal reference images into a memory bank (Roth et al., 2022) and calculates the anomaly score with the cosine similarity between the query image and its most similar image in the memory bank.
@@ 5.2.2 ID Training-based Methods
Training-based methods for VLM-based AD are generally categorized as auxiliary training (in Sec. 5.1.2), but some approaches use data from the same domain as the test data for training. CLIP-FSAC (Zuo et al., 2024) enhances few-shot anomaly classification by introducing a two-stage fine-tuning framework with modality-specific adapters, where the image adapter is trained first using text features, followed by the text adapter while freezing the image adapter. To address data scarcity, it employs synthetic anomaly generation using random perturbation and Poisson-based editing to create diverse training samples. PromptAD (Li et al., 2024c) proposes a prompt learning method for one-class AD (where the normal class consists of one class). In one-class AD, traditional prompt learning methods for multi-class classification (e.g., CoOp (Zhou et al., 2022c)) do not work well. To address this, PromptAD creates a large number of anomaly prompts by adding a learnable anomaly suffix to the normal prompt. It then learns to bring the visual features closer to the normal prompt and further away from the anomaly prompts. As a different direction, Kwak et al. (2024) proposed ADP, which learns anomaly-aware prompts using the personalization capabilities of diffusion models. Also, Li et al. (2024f) has proposed a novel anomaly personalization approach, which applies a one-to-normal transformation to query images using a customized anomaly-free generation model, ensuring alignment with the normal data manifold. This method leverages a triplet contrastive anomaly inference strategy, incorporating comparisons between the query image, a generated anomaly-free data pool, and textual prompts to enhance prediction stability and robustness.
@@ 5.2.3 Auxiliary Training- and Reference-based Methods
We explore the methods trained on auxiliary datasets and utilize the normal images in the target domain as references during inference. An early work in this category is APRIL-GAN (few-shot) (Chen et al., 2023b), which uses a linear layer trained with auxiliary datasets. Similar to WinCLIP+ (Jeong et al., 2023), APRIL-GAN (few-shot) utilizes a few ID reference images with a memory bank-based approach (Roth et al., 2022). More recently, Zhu & Pang (2024) proposed an in-context-learning-based method called InCTRL. InCTRL trains a model to discriminate anomalies from normal samples by learning to identify residuals or discrepancies between query images and a set of few-shot normal images (in-context sample prompts) from auxiliary data. During inference, InCTRL identifies anomalies by measuring the discrepancy between the features of the query image and a few in-context normal samples from the target dataset.
@@ 5.3.1 Anomaly Detection with Localization Models
Some works (Cao et al., 2023a; Li et al., 2024a) tackle AD using foundation models for localization, such as SAM (Kirillov et al., 2023) or GroundingDINO (Liu et al., 2024e). SAA (Cao et al., 2023a) is a pioneering approach that integrates SAM into anomaly segmentation tasks. It first employs prompt-driven object detection techniques such as GroundingDINO (Liu et al., 2024e) to generate box-level regions conditioned on prompts, highlighting the target anomaly areas. These generated boxes serve as prompts for SAM, which then produces the final anomaly segmentation results. Building on this, SAA+ (Cao et al., 2023a) introduces hybrid prompts that blend domain-specific knowledge with contextual image information, helping to reduce ambiguities inherent in language-based prompts. In contrast, CLIPSAM (Li et al., 2025) replaces GroundingDINO with CLIP, leveraging its superior localization capability to provide more precise prompts for SAM. To bridge the domain gap from natural images,  Yang et al. (2025) propose a promptable anomaly segmentation model with SAM, incorporating a novel Self-Perception Tuning (SPT) method. They first design a Self-Draft Tuning strategy, in which SAM initially generates a coarse draft of the anomaly mask, followed by a mask refinement process. Furthermore, they introduce a Visual-Relation-Aware Adapter to enhance the internal perception knowledge of discriminative relation information during decoding.
Given that AD necessitates localization, it is expected that the number of works employing localization foundation models such as SAM will continue to increase.
@@ 5.3.2 Medical Anomaly Detection
While most works on VLM-based AD focus on industrial AD, recent studies have begun to challenge medical anomaly detection (medical AD) (Chen et al., 2023c; Zhu & Pang, 2024; Huang et al., 2024; Hua et al., 2024; Zhang et al., 2024c; Cao et al., 2024c). VLM-based medical AD is a challenging area as well as industrial AD due to the larger gap between different data modalities. A representative work on medical AD is MVFA (Huang et al., 2024). MVFA is a method specifically tailored for medical AD. It incorporates multiple residual adapters into the CLIP’s visual encoder to reduce the domain gap, enabling a stepwise enhancement of visual features across different levels. The future progression of medical AD and industrial AD offers an intriguing perspective, exploring whether these fields will develop independently or influence each other. However, when considering practical applications, it should be noted that medical AD faces the challenge that anomalies are not always describable in the language. Therefore, the development of medical AD methods that do not use CLIP is also important.
5.3.3 Video Anomaly Detection
Compared to OOD detection, one important point is that anomaly detection has been actively studied not only in image domains but also in video domains (Zhu et al., 2024b). Video Anomaly Detection (VAD) aims to automatically detect unusual events in videos, and it has many applications such as surveillance and monitoring (Wang et al., 2019b). VAD has several challenges. First, there is no clear and unified definition of anomalies, because what is normal or abnormal (for example, walking on a sidewalk vs. on a highway) depends on the context. Second, anomalies happen rarely and unpredictably, so it is difficult to collect high-quality datasets. This makes it hard to learn such patterns effectively. Because of these challenges and interests, VAD has become an active field, and many survey papers have been published (Suarez & Naval Jr, 2020; Abbas & Al-Ani, 2022; Jiao et al., 2023; Zhu et al., 2024b; Abdalla et al., 2024).
Before the VLM era, most VAD research focused on building task-specific models using unsupervised learning, one-class learning, or weakly supervised learning, because there was a lack of labeled anomaly data (Wu et al., 2024a). After the introduction of VLMs, the research direction changed a lot. Using powerful features from pre-trained VLMs, a lot of work tried zero-shot and few-shot VAD with much less task-specific data (Zanella et al., 2024b). Furthermore, the language understanding capabilities of VLMs have enabled new task settings, such as the integration of textual information (Zanella et al., 2024a; b) and the generation of natural language explanations for detected anomalies (Ye et al., 2025; Zhang et al., 2025a).
These recent trends show that the research direction of VAD has changed a lot with VLMs. On the other hand, OOD detection in video domains is still not widely studied. In this survey, our main goal is to better understand OOD detection itself, while referring to related areas such as AD only as supplementary information. Therefore, we do not give a full review of VAD in this survey, and instead refer readers to existing VAD survey papers (Suarez & Naval Jr, 2020; Abbas & Al-Ani, 2022; Jiao et al., 2023; Zhu et al., 2024b; Abdalla et al., 2024).
@@ 5.4 Discussion
We discuss the similarities and differences between VLM-based OOD detection and VLM-based AD to deepen our understanding of VLM-based OOD detection.
@@ 5.4.1 Difference between Each Methodology
Differing Scopes of OOD  OOD detection and AD differ significantly in the scope of OOD (anomaly) they cover, which leads to differences in methodologies, particularly in the use of OOD prompts. As explained in Sec. 3, sensory AD is intended for specific use cases like industrial inspection, where samples deviating from predefined normality (e.g., defective products) are considered anomalies (Yang et al., 2024; Ruff et al., 2021). In other words, in sensory AD, the anomaly space is limited to damaged objects with shared semantics, and anomalies like images of dogs are not expected. This limited anomaly space allows even simple prompts to achieve decent performance. Therefore, as shown in Table 2, all AD methods utilize anomaly prompts. Conversely, in OOD detection, as explained in Sec. 3.1, anything semantically different from the ID class is considered OOD. Thus, utilizing naive manual OOD prompts is forbidden (even if it improves benchmark performance). This vastness of the OOD space is the key factor differentiating the methodologies between the two fields.
Difficulty of Localization Task  VLM-based OOD detection and VLM-based AD differ significantly in the difficulty of OOD (anomaly) localization tasks. In VLM-based AD, anomaly segmentation is a mainstream task, performed alongside classification in many papers. However, in VLM-based OOD detection, there has been no research on object-level OOD detection/segmentation. Object-level OOD detection aims to detect OOD objects (Du et al., 2022c; b; a). The inactivity is related to the size of the OOD space, and the too-vast space of OOD makes it difficult to identify OOD objects with prompts effectively. To pave the way for future development, the foundation models for localization such as SAM (Kirillov et al., 2023), which can segment individual objects, have the potential to address object-level OOD detection/segmentation. Object-level OOD detection/segmentation using SAM is a promising future research direction.